# Learning to Compress Prompt in Natural Language Formats

### v1

@@ 1 Introduction
Figure 1: An example of successful prompt compression with NL formats. The compressed NL-formatted prompt (green) aims to obtain a shorter length and maintain transferability and utility simultaneously than the original long prompt (red).
@@ 4.1 Dataset
Claude2 76.6 69.4 74.6 85.6 52.7 84.9 59.4 58.2 95.0 90.1
Length (# of Token) 831 154 751 231 378.39 95.66 915.7 171.61
Compress Rate (%) 81.4% 69.3% 74.71% 81.25%
@@ 4.3 Main Results (RQ1)
In the primary experiment, we establish a compression constraint limiting to a maximum of 150 tokens for the CSQA, MultiRC, and TriviaQA-Long datasets and a maximum of 300 tokens for the GSM8K dataset, where 150 tokens are not sufficient for preserving the logic of GSM8K dataset. We observe that the Nano-Capsulator obtains up to 81.4% of the compression rate and saves up to 80.1% of the Claude2 API cost compared to the original input prompts, as displayed in Table 2. The cost of PaLM API can be found in Appendix C. In terms of utility preservation, Nano-Capsulator retains the original performance on the CSQA and GSM8k datasets mostly among three LLMs. Remarkably, Nano-Capsulator maintains a performance level almost identical to that achieved with non-compressed prompts in MultiRC datasets. The significant compression rate achieved can advantageously impact the model by allowing for a higher tolerance in batch inference, accompanied by reduced latency and cost.
@@ Appendix F Instruction Usage in Inference LLMs
Compression Tasks
$\textbf{T}_{\text{Rep}}$ (Replicating Instruction)
$\textbf{T}_{\text{Summ}}$ (Summarizing Instruction)
Few-shot CoT
Repeat the following main input.
Please summarize each question-answer pair in one sentence within less than {word count} words. Make sure not to repeat the input question-answer pair.
Reading Comprehension
Repeat the following main input.
Please summarize the passage within less than {word count} words. Make sure not to repeat the passage.

### v2

@@ 1 Introduction
Figure 1: An example of successful prompt compression with NL formats. The compressed NL-formatted prompt (green) aims to obtain a shorter length and maintain transferability and utility of the long prompt (red).
@@ 4.1 Dataset
Claude2 76.6 69.4 74.6 85.6 52.7 84.9 59.4 58.2 95.0 92.3
Length (# of Token) 831 154 751 231 378.39 95.66 915.7 422.6
Compress Rate (%) 81.4% 69.3% 74.71% 53.84%
@@ 4.3 Main Results (RQ1)
In the primary experiment, we establish a compression constraint limiting to a maximum of 150 tokens for the CSQA and MultiRC datasets; and a maximum of 350 and 500 tokens for the GSM8K and TriviaQA-Long dataset, where 150 tokens are not sufficient for preserving the logic of GSM8K and TriviaQA-Long dataset. We observe that the Nano-Capsulator obtains up to 81.4% of the compression rate and saves up to 80.1% of the Claude2 API cost compared to the original input prompts, as displayed in Table 2. The cost of PaLM API can be found in Appendix C. For utility preservation, Nano-Capsulator retains the original performance on the CSQA, GSM8k, and TriviaQA-Long datasets mostly among three LLMs. Remarkably, Nano-Capsulator maintains almost identical performance to that achieved with non-compressed prompts in MultiRC datasets. The significant compression rate can advantageously impact the LLMs by allowing for a higher tolerance in batch inference, accompanied by reduced latency and cost.
@@ Appendix F Instruction Usage in Inference LLMs
Compression Tasks $\textbf{T}_{\text{Rep}}$ (Replicating Instruction) $\textbf{T}_{\text{Summ}}$ (Summarizing Instruction)
Few-shot CoT Repeat the following main input. Please summarize each question-answer pair in one sentence within less than {word count} words. Make sure not to repeat the input question-answer pair.
Reading Comprehension Repeat the following main input. Please summarize the passage within less than {word count} words. Make sure not to repeat the passage.